OpenStudy (buttermequeasy):

[4_07]Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6. y = 3x - 10 y = 3x - 6 y = 3x - 2 y = 3x - 12 Urgent help needed!

OpenStudy (buttermequeasy):

How do you exactly do this?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

y+4=3(x-2) y+4=3x-6 y=3x-10

OpenStudy (anonymous):

hope this helps :)

OpenStudy (buttermequeasy):

I see. Thank you!

OpenStudy (anonymous):

anytime :)